Mathematics 133 – Intensive Calculus for Science 1 Quiz 1 – September 14, 2001

Your Name:

Do all work on this sheet (front and back if needed). You have 20 minutes.

A) The following is the graph y = f(x). On the same set of axes, sketch the graph y = −3f(x + 1).

B) Let f(x) = x^{2} − 2x and g(x) = x + 1. Find and simplify the composition f(g(x)).

C) Daily sales S (in dollars) at “Cool Beans 2” (the coffee kiosk outside the Science Library downstairs) were S = 200e^{(.001)t} on day t after it opened. Will sales ever reach $500? On what day? (Leave your answer in “exact” form.)